# GUI‑AIMA: Aligning Intrinsic Multimodal Attention with a Context Anchor for GUI Grounding

### v1

@@ 4 Experiments
Model CAD Dev Creative Scientific Office OS Average Text Icon Text Icon Text Icon Text Icon Text Icon Text Icon Text Icon Avg. General GPT-4o 2.0 0.0 1.3 0.0 1.0 0.0 2.1 0.0 1.1 0.0 0.0 0.0 1.3 0.0 0.8 Claude Computer 14.5 3.7 22.0 3.9 25.9 3.4 33.9 15.8 30.1 16.3 11.0 4.5 23.4 7.1 17.1 Qwen2.5-VL-3B 9.1 7.3 22.1 1.4 26.8 2.1 38.2 7.3 33.9 15.1 10.3 1.1 23.6 3.8 16.1 Qwen2.5-VL-7B 16.8 1.6 46.8 4.1 35.9 7.7 49.3 7.3 52.5 20.8 37.4 6.7 38.9 7.1 26.8 SFT OS-Atlas-7B 12.2 4.7 33.1 1.4 28.8 2.8 37.5 7.3 33.9 5.7 27.1 4.5 28.1 4.0 18.9 UGround-V1-7B 15.8 1.2 51.9 2.8 47.5 9.7 57.6 14.5 60.5 13.2 38.3 7.9 45.2 8.1 31.1 UI-TARS-72B 18.8 12.5 62.9 17.2 57.1 15.4 64.6 20.9 63.3 26.4 42.1 15.7 50.9 17.6 38.1 Jedi-3B 27.4 9.4 61.0 13.8 53.5 8.4 54.2 18.2 64.4 32.1 38.3 9.0 49.8 13.7 36.1 Jedi-7B 38.0 14.1 42.9 11.0 50.0 11.9 72.9 25.5 75.1 47.2 33.6 16.9 52.6 18.2 39.5 UI-TARS-1.5-7B 38.6 11.0 58.4 12.4 58.1 15.4 66.7 21.9 74.6 35.9 49.5 13.5 57.5 16.9 42.0 GUI-Actor-3B - - - - - - - - - - - - - - 42.2 GUI-Actor-7B - - - - - - - - - - - - - - 44.6 RL UI-R1-E-3B 37.1 12.5 46.1 6.9 41.9 4.2 56.9 21.8 65.0 26.4 32.7 10.1 - - 33.5 InfiGUI-R1-3B 33.0 14.1 51.3 12.4 44.9 7.0 58.3 20.0 65.5 28.3 43.9 12.4 49.1 14.1 35.7 GUI-G1-3B 39.6 9.4 50.7 10.3 36.6 11.9 61.8 30.0 67.2 32.1 23.5 10.6 49.5 16.8 37.1 SE-GUI-3B 38.1 12.5 55.8 7.6 47.0 4.9 61.8 16.4 59.9 24.5 40.2 12.4 50.4 11.8 35.9 GUI-G2-3B 20.3 9.4 42.2 2.1 43.4 2.8 43.8 10.0 46.3 15.1 29.0 4.5 37.6 6.0 25.5 GUI-AIMA-3B 43.7 25.0 63.0 20.0 57.1 17.5 68.1 33.6 69.5 35.9 47.7 22.5 58.1 24.2 45.2 GUI-AIMA-3B $(soft)^{*}$ 53.3 23.4 63.0 24.1 55.1 14.7 66.7 30.9 76.8 34.0 55.1 21.4 61.6 23.5 47.1 GUI-AIMA-3B + zoom-in 66.5 34.4 75.3 33.1 64.7 23.1 75.7 38.2 85.9 62.3 65.4 49.4 72.1 36.8 58.6
@@
We apply Qwen2.5-VL-3B-Instruct (Bai et al., 2025) as the MLLM backbone for GUI-AIMA and all ablation variants. We follow the special token setting in GUI-Actor (Wu et al., 2025) and retain the next-token prediction loss for GUI-AIMA’s grounding format. The training of GUI-AIMA-3B is conducted on 8 NVIDIA A100-80G GPUs with effective global batch size 64 and learning rate 5e-6. We set $\alpha$ as 0.8 for the adaptive deviation control in Eq. 2. Regarding the setting of visual-sink query token $\mathcal{Q}_{s}$, we select the top-1 global large tokens complying with Eq. 9. As for the training dataset, we employ the entire training set of GUIAct (Chen et al., 2024a), AndroidControl (Li et al., 2024), Wave-UI (Jeffries & Team, 2024), randomly sampled 60k samples from UGround (Gou et al., 2024) and from GTA1 training set (Yang et al., 2025), respectively, with 259k instructions and $\sim$85k images in all. For the 45k ablation training set for Section 4.2, we select the first 10k samples from GUIAct, AndroidControl, Wave-UI and first 15k samples from UGround. GUI-AIMA-3B is trained for one epoch with all parameters unfrozen, without extra modules and a warmup stage. Implementation of baselines is in Appendix B. For two-step inference introduced in Section 3.5, we set the crop size to $448$ pixels and adopt a $2\times$ zoom-in ratio for the second pass.
@@
We evaluate GUI-AIMA, baselines and ablation variants on ScreenSpot-v2 (Wu et al., 2024) for general GUI visual grounding evaluation across mobile, desktop and web domains, ScreenSpot-Pro (Li et al., 2025) for testing on challenging higher-resolution screenshots from diverse and complex professional software scenarios, e.g. interfaces from different Operating Systems with large distribution gap, and OSWorld-G (Xie et al., 2025a). All benchmarks contain separated text-centered and icon-centered visual grounding tasks, where the icon set is the more abstract visual-based task less hinted by text. For grounding accuracy, we follow the standard center point-based metric (Lin et al., 2024) that the correct click-point prediction locates within the ground-truth bounding box.
@@
Three categories of methods are compared: (1) General MLLMs including GPT-4o (OpenAI, 2024), Gemini-2.5-Pro (Comanici et al., 2025), Claude Computer (Hu et al., 2024), Operator (OpenAI, 2025), Qwen2.5-VL (Bai et al., 2025); (2) GUI-specific supervised fine-tuned model: OS-Atlas (Wu et al., 2024), UGround (Gou et al., 2024), UI-TARS (Qin et al., 2025), JEDI (Xie et al., 2025b) and GUI-Actor (Wu et al., 2025) ; (3) GUI-specific Reinforcement fine-tuned model: UI-R1 (Lu et al., 2025), InfiGUI-R1 (Liu et al., 2025), GUI-G1 (Zhou et al., 2025), SE-GUI (Yuan et al., 2025), GUI-G2 (Tang et al., 2025).
@@ 4.1 Main Results: Grounding Performance
We compare GUI-AIMA with state-of-the-art methods on ScreenSpot-Pro, ScreenSpot-v2 and OSWorld-G and report results in Table 1, Table 2 and Table 3. While only being trained with 259k samples without filtering or careful selection, GUI-AIMA-3B outperforms all same scale 3B MLLM-based models and shows comparable or even better results than larger scaled models. Among Qwen-2.5-VL based models, marked with in both tables, GUI-AIMA-3B can better handle the visual grounding in complex software scenarios across different platforms and achieves the best performance on the most challenging ScreenSpot-Pro benchmark, especially on the abstract icon task set. Specifically, GUI-AIMA-3B is better than strong large size coordinate-based UI-TARS-1.5-7B, JEDI-7B, and also better than the embedding-based coordinate-free GUI-Actor-7B model, highlighting the superiority of directly supervising on the multi-head self-attention weights instead of modeling the query-visual attention map via hidden states. On ScreenSpot-v2, GUI-AIMA-3B achieve the comparable results with the strongest baselines, such as JEDI-7B and UI-TARS-7B, and better than the same size GUI-Actor-3B. Besides the advanced performance, another advantage is the training efficiency of GUI-AIMA with only 259k training elements, as most supervised fine-tuned baselines trained on millions of GUI elements. For the comparison with reinforcement fine-tuned baselines, while both trained with smaller training sets than SFT baselines, GUI-AIMA-3B performs better and shows better generalization. Among GUI-AIMA-3B and GUI-AIMA-3B (soft), GUI-AIMA-3B (soft) handles the high-resolution interfaces in ScreenSpot-Pro with 1.9% improvments, and GUI-AIMA-3B is better on dealing with diverse graphic environments in ScreenSpot-v2 and OSWorld-G. The two-step inference with zoom-in significantly improve the performance of GUI-AIMA on high-resolution benchmarks, ScreenSpot-pro and OSWorld-G, demonstrating the flexibility of GUI-AIMA’s attention-based patch-wise grounding for inference-time improvements.
@@ 4.2 Ablations
Here, we also justify the overlapping- and distance-aware (Tang et al., 2025) patch-wise labeling for fine-tuning GUI-AIMA in Eq. 2. Based on “global $\operatorname{top-1}\mathcal{Q}_{s}$”, down-weighting partial positive and distanced image patch instead of labeling all patch equally leads to 1.26% further enhancement on ScreenSpot-Pro.
@@ Appendix A Analysis for Visual-sink Query Token
The observations above verify the claim in paper that the query-visual pattern discovered in hidden states $\mathbf{H}^{l}$ is not necessarily statistically prevailing among each head’s self-attention matrix, as only a smaller subset of attention heads are “semantic heads” that key on semantic functionality and representation similarity, which is also supported by Elhelo & Geva (2024); Olsson et al. (2022); Voita et al. (2019).

### v2

@@ 4 Experiments
Model CAD Dev Creative Scientific Office OS Average Text Icon Text Icon Text Icon Text Icon Text Icon Text Icon Text Icon Avg. General GPT-4o 2.0 0.0 1.3 0.0 1.0 0.0 2.1 0.0 1.1 0.0 0.0 0.0 1.3 0.0 0.8 Claude Computer 14.5 3.7 22.0 3.9 25.9 3.4 33.9 15.8 30.1 16.3 11.0 4.5 23.4 7.1 17.1 Qwen2.5-VL-3B 9.1 7.3 22.1 1.4 26.8 2.1 38.2 7.3 33.9 15.1 10.3 1.1 23.6 3.8 16.1 Qwen2.5-VL-7B 16.8 1.6 46.8 4.1 35.9 7.7 49.3 7.3 52.5 20.8 37.4 6.7 38.9 7.1 26.8 SFT OS-Atlas-7B 12.2 4.7 33.1 1.4 28.8 2.8 37.5 7.3 33.9 5.7 27.1 4.5 28.1 4.0 18.9 UGround-V1-7B 15.8 1.2 51.9 2.8 47.5 9.7 57.6 14.5 60.5 13.2 38.3 7.9 45.2 8.1 31.1 UI-TARS-72B 18.8 12.5 62.9 17.2 57.1 15.4 64.6 20.9 63.3 26.4 42.1 15.7 50.9 17.6 38.1 Jedi-3B 27.4 9.4 61.0 13.8 53.5 8.4 54.2 18.2 64.4 32.1 38.3 9.0 49.8 13.7 36.1 Jedi-7B 38.0 14.1 42.9 11.0 50.0 11.9 72.9 25.5 75.1 47.2 33.6 16.9 52.6 18.2 39.5 UI-TARS-1.5-7B 38.6 11.0 58.4 12.4 58.1 15.4 66.7 21.9 74.6 35.9 49.5 13.5 57.5 16.9 42.0 GUI-Actor-3B - - - - - - - - - - - - - - 42.2 GUI-Actor-7B - - - - - - - - - - - - - - 44.6 RL UI-R1-E-3B 37.1 12.5 46.1 6.9 41.9 4.2 56.9 21.8 65.0 26.4 32.7 10.1 - - 33.5 InfiGUI-R1-3B 33.0 14.1 51.3 12.4 44.9 7.0 58.3 20.0 65.5 28.3 43.9 12.4 49.1 14.1 35.7 GUI-G1-3B 39.6 9.4 50.7 10.3 36.6 11.9 61.8 30.0 67.2 32.1 23.5 10.6 49.5 16.8 37.1 SE-GUI-3B 38.1 12.5 55.8 7.6 47.0 4.9 61.8 16.4 59.9 24.5 40.2 12.4 50.4 11.8 35.9 GUI-G2-3B 20.3 9.4 42.2 2.1 43.4 2.8 43.8 10.0 46.3 15.1 29.0 4.5 37.6 6.0 25.5 GUI-AIMA-3B 44.7 21.9 68.2 28.3 59.6 21.7 69.4 37.3 70.6 49.1 65.4 31.5 62.0 30.0 49.8 GUI-AIMA-3B $(soft)^{*}$ 53.3 15.6 62.3 26.9 58.1 15.4 72.2 35.5 73.5 49.1 62.6 30.3 63.2 27.0 49.3 GUI-AIMA-3B + zoom-in 65.0 34.4 72.7 36.6 66.7 23.8 81.3 41.8 85.3 60.4 69.2 46.1 73.1 37.8 59.6
@@
We apply Qwen2.5-VL-3B-Instruct (bai2025qwen2) as the MLLM backbone for GUI-AIMA and all ablation variants. We follow the special token setting in GUI-Actor (wu2025gui) and retain the next-token prediction loss for GUI-AIMA’s grounding format. The training of GUI-AIMA-3B is conducted on 8 NVIDIA A100-80G GPUs with effective global batch size 64 and learning rate 5e-6. We set $\alpha$ as 0.8 for the adaptive deviation control in Eq. 2. Regarding the setting of visual-sink query token $\mathcal{Q}_{s}$, we select the top-1 global large tokens complying with Eq. 9. As for the training dataset, we employ the entire training set of GUIAct (chen2024guicourse), AndroidControl (li2024effects), Wave-UI (waveui), randomly sampled 60k samples from UGround (gou2024navigating) and from GTA1 training set (yang2025gta1guitesttimescaling), respectively, with 259k instructions and $\sim$85k images in all. For the 45k ablation training set for Section 4.2, we select the first 10k samples from GUIAct, AndroidControl, Wave-UI and first 15k samples from UGround. GUI-AIMA-3B is trained for one epoch with all parameters unfrozen, without extra modules and a warmup stage. Implementation of baselines is in Appendix B. For two-step inference introduced in Section 3.5, we set the crop size to $448$ pixels and adopt a $2\times$ zoom-in ratio for the second pass.
Efficient Extraction of Self-Attention Map
Fast and memory-efficient attention implementations, such as FlashAttention dao2022flashattention, avoid materializing the full attention matrix and therefore do not store intermediate attention weights. By contrast, the eager execution in original transformer attention vaswani2017attention can return the full attention maps, but it is neither fast nor memory-efficient, which becomes a bottleneck for large-scale training and for scaling up model size. In GUI-AIMA, we address this by combining FlashAttention with eager attention: we use FlashAttention for the regular layer-to-layer forward pass, and then reuse the same attention parameters to compute a partial attention map, covering the rows of text tokens and the <ANCHOR> token, via eager execution for GUI-AIMA’s attention grounding. Since visual tokens dominate the attention in MLLMs, the additional computation and memory cost is ignorable.
@@
We evaluate GUI-AIMA, baselines and ablation variants on ScreenSpot-v2 (wu2024atlas) for general GUI visual grounding evaluation across mobile, desktop and web domains, ScreenSpot-Pro (li2025screenspot) for testing on challenging higher-resolution screenshots from diverse and complex professional software scenarios, e.g. interfaces from different Operating Systems with large distribution gap, and OSWorld-G (xie2505scaling). All benchmarks contain separated text-centered and icon-centered visual grounding tasks, where the icon set is the more abstract visual-based task less hinted by text. For grounding accuracy, we follow the standard center point-based metric (lin2024showui) that the correct click-point prediction locates within the ground-truth bounding box.
@@
Three categories of methods are compared: (1) General MLLMs including GPT-4o (openai2024gpt4o), Gemini-2.5-Pro (comanici2025gemini), Claude Computer (hu2024dawnguiagentpreliminary), Operator (cua2025), Qwen2.5-VL (bai2025qwen2); (2) GUI-specific supervised fine-tuned model: OS-Atlas (wu2024atlas), UGround (gou2024navigating), UI-TARS (qin2025ui), JEDI (xie2025scaling) and GUI-Actor (wu2025gui) ; (3) GUI-specific Reinforcement fine-tuned model: UI-R1 (lu2025ui), InfiGUI-R1 (liu2025infiguig1advancingguigrounding), GUI-G1 (zhou2025gui), SE-GUI (yuan2025enhancing), GUI-G2 (tang2025gui).
@@ 4.1 Main Results: Grounding Performance
We compare GUI-AIMA with state-of-the-art methods on ScreenSpot-Pro, ScreenSpot-v2 and OSWorld-G and report results in Table 1, Table 2 and Table 3. While only being trained with 259k samples without filtering or careful selection, GUI-AIMA-3B outperforms all same scale 3B MLLM-based models and shows comparable or even better results than larger scaled models. Among Qwen-2.5-VL based models, marked with in both tables, GUI-AIMA-3B can better handle the visual grounding in complex software scenarios across different platforms and achieves the best performance on the most challenging ScreenSpot-Pro benchmark, especially on the abstract icon task set. Specifically, GUI-AIMA-3B is better than strong large size coordinate-based UI-TARS-1.5-7B, JEDI-7B, and also better than the embedding-based coordinate-free GUI-Actor-7B model, highlighting the superiority of directly supervising on the multi-head self-attention weights instead of modeling the query-visual attention map via hidden states. On ScreenSpot-v2, GUI-AIMA-3B achieve the comparable results with the strongest baselines, such as JEDI-7B and UI-TARS-7B, and better than the same size GUI-Actor-3B, while GUI-AIMA-3B is trained with much less web data. Besides the advanced performance, another advantage is the training efficiency of GUI-AIMA with only 259k training elements, as most supervised fine-tuned baselines trained on millions of GUI elements. For the comparison with reinforcement fine-tuned baselines, while both trained with smaller training sets than SFT baselines, GUI-AIMA-3B performs better and shows better generalization. Among GUI-AIMA-3B and GUI-AIMA-3B (soft), GUI-AIMA-3B is slightly better on dealing with diverse graphic environments in ScreenSpot-v2, OSWorld-G and ScreenSpot-pro. The two-step inference with zoom-in without extra training significantly improves the performance of GUI-AIMA on high-resolution benchmarks, ScreenSpot-pro and OSWorld-G, specifically 59.6% on ScreenSpot-pro using GUI-AIMA-3B (soft) and 63.8% on OSWorld-G using GUI-AIMA-3B, demonstrating the flexibility of GUI-AIMA’s attention-based patch-wise grounding for inference-time improvements.
@@ 4.2 Ablations
Here, we also justify the overlapping- and distance-aware (tang2025gui) patch-wise labeling for fine-tuning GUI-AIMA in Eq. 2. Based on “global $\operatorname{top-1}\mathcal{Q}_{s}$”, down-weighting partial positive and distanced image patch instead of labeling all patch equally leads to 1.26% further enhancement on ScreenSpot-Pro.
@@ Appendix A Analysis for Visual-sink Query Token
The observations above verify the claim in paper that the query-visual pattern discovered in hidden states $\mathbf{H}^{l}$ is not necessarily statistically prevailing among each head’s self-attention matrix, as only a smaller subset of attention heads are “semantic heads” that key on semantic functionality and representation similarity, which is also supported by elhelo2024inferring; olsson2022context; voita2019analyzing.